A firm is employing 100 workers at a wage rate(W) = $l5lhour and 50 units of capital at a


rental rate® = $30lhour. At the firm's current input use, the marginal product of labor is 45 and


the marginal product of capital is 60. Is the firm producing its current level of output at the


minimum cost


a) Yes because MPL/W > MPK/R.


b) Yes because W<R


c) No because MPL/W > MPK/R


d) No because MPL < MPK

1
Expert's answer
2022-01-19T13:13:00-0500

"\frac{100}{15}\cdot 45>\frac{100}{30}\cdot 60,"


answer is c) No because MPL/W > MPK/R.


for the firm to produce its current level of output at the minimum cost

MPL / w = MPK / r


interpretation

1 / MPL more labor at cost w / MPL

Or 1 / MPK more capital at cost r / MPK

or a combination of the two. These two must be equal when making the best decision.